# FedIA: A Plug-and-Play Importance-Aware Gradient Pruning Aggregation Method for Domain-Robust Federated Graph Learning on Node Classification

### v1

@@ Abstract
Federated Graph Learning (FGL) under domain skew—as observed on platforms such as Twitch Gamers and multilingual Wikipedia networks—drives client models toward incompatible representations, rendering naïve aggregation both unstable and ineffective. We find that the culprit is not the weighting scheme but the noisy gradient signal: empirical analysis of baseline methods suggests that a vast majority of gradient dimensions can be dominated by domain-specific variance. We therefore shift focus from ”aggregation-first” to a projection-first strategy that denoises client updates before they are combined. The proposed FedIA framework realises this Importance-Aware idea through a two-stage, plug-and-play pipeline: (i) a server-side top-$\rho$ mask keeps only the most informative $\sim 5\%$ of coordinates, and (ii) a lightweight influence-regularised momentum weight suppresses outlier clients. FedIA adds no extra uplink traffic and only negligible server memory, making it readily deployable. On both homogeneous (Twitch Gamers) and heterogeneous (Wikipedia) graphs, it yields smoother, more stable convergence and higher final accuracy than nine strong baselines. A convergence sketch further shows that dynamic projection maintains the optimal $\mathcal{O}\!\bigl{(}\sigma^{2}/\sqrt{T}\bigr{)}$ rate. CODE: https://github.com/itoritsu/FedIA.
@@
Prior solutions primarily fall into two impractical categories. Parameter regularization methods (e.g., FedProx (Li et al. 2020), FedDyn (Acar et al. 2021)) attempt to penalize deviations from global parameters but fail to explicitly handle gradient-level noise. Prototype-based or mask-based methods (e.g., FGGP (Wan, Huang, and Ye 2024), FedHEAL (Chen, Huang, and Ye 2024)) attempt to synchronize domain representations by exchanging prototypes or maintaining historical client masks, but incur prohibitive communication and memory overheads, respectively. Critically, none of these methods tackles the root cause—gradient noise—prior to aggregation.
We argue that aggregation is not the primary issue to solve; pre-conditioning the gradient signals is. Our analysis on baseline models like FedAvg reveals that a significant portion of gradient coordinates, potentially over 90% in some cases, can carry domain-specific noise, and retaining a small subset (around 5%) suffices to maintain robust convergence. Moreover, we emphasize that directly transmitting explicit gradients—rather than weight updates—eliminates server-side inference noise, further ensuring high-quality gradient signals for aggregation. This insight motivates a fundamental paradigm shift toward a projection-dominant approach, prioritizing gradient filtering before aggregation.
@@ 2.2 Related Works
Federated Graph Learning Approaches. FGL combines FL and GNN to leverage distributed graph data while preserving privacy. Existing methods broadly include FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), FedOPT (Reddi et al. 2020), FedDyn (Acar et al. 2021), and MOON (Li, He, and Song 2021), originally developed for image classification, which directly aggregate parameters or rely on model regularization to enhance stability. However, these parameter-based approaches inherently lack mechanisms to resolve domain-induced discrepancies explicitly, thus exhibiting degraded performance under severe graph domain skew.
@@ 3 Challenges and Insights
Representation divergence manifests as high-variance, often conflicting gradients. Baseline FedAvg oscillates on a rugged loss surface (Fig. 2(b)), while a state-of-the-art aggregator, FGGP, remains unstable across settings (Fig. 2(c)). Simply exchanging larger messages—prototypes, local statistics or second moments—does not tame the noise; the optimiser continues to chase contradictory directions.

### v2

@@ Abstract
Federated Graph Learning (FGL) under domain skew drives client models toward incompatible representations, rendering naïve aggregation both unstable and ineffective. We find that the culprit is not the weighting scheme but the noisy gradient signal: empirical analysis of baseline methods suggests that a vast majority of gradient dimensions can be dominated by domain-specific variance. We therefore shift focus from ”aggregation-first” to a projection-first strategy that denoises client updates before they are combined. The proposed FedIA framework realises this Importance-Aware idea through a two-stage, plug-and-play pipeline: (i) a server-side top-$\rho$ global importance mask can even keep only the most informative $\sim 5\%$ of coordinates, and (ii) a lightweight influence-regularised momentum weight suppresses outlier clients. FedIA adds no extra uplink traffic and only negligible server memory, making it readily deployable. On both homogeneous (Twitch Gamers) and heterogeneous (Wikipedia) graphs, it yields smoother, more stable convergence and higher final accuracy than nine strong baselines. A convergence sketch further shows that dynamic projection maintains the optimal $\mathcal{O}\!\bigl(\sigma^{2}/\sqrt{T}\bigr)$ rate. CODE: https://github.com/itoritsu/FedIA.
@@
Prior solutions primarily fall into two impractical categories. Parameter regularization methods (e.g., FedProx (Li et al. 2020), FedDyn (Acar et al. 2021)) attempt to penalize deviations from global parameters but fail to explicitly handle gradient-level noise. Prototype-based or mask-based methods (e.g., FGGP (Wan, Huang, and Ye 2024), FedHEAL (Chen, Huang, and Ye 2024)) attempt to synchronize domain representations by exchanging prototypes or maintaining historical client masks, but incur prohibitive communication and memory overheads, respectively. Critically, none of these methods tackles the root cause by gradient noise, prior to aggregation.
We argue that aggregation is not the primary issue to solve; pre-conditioning the gradient signals is. Our analysis on baseline models like FedAvg reveals that a significant portion of gradient coordinates, potentially over 90% in some cases, can carry domain-specific noise, and retaining a small subset (around 5%) suffices to maintain robust convergence. Moreover, we emphasize that directly transmitting explicit gradients eliminates server-side inference noise, rather than weight updates, further ensuring high-quality gradient signals for aggregation. This insight motivates a fundamental paradigm shift toward a projection-dominant approach, prioritizing gradient filtering before aggregation.
@@ 2.2 Related Works
From FL’s Heterogeneity Solutions to the FGL’s. FGL combines FL and GNN to leverage distributed graph data while preserving privacy. Existing methods broadly include FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), FedOPT (Reddi et al. 2020), FedDyn (Acar et al. 2021), and MOON (Li, He, and Song 2021), originally developed for image classification, which directly aggregate parameters or rely on model regularization to enhance stability. However, these parameter-based approaches inherently lack mechanisms to resolve domain-induced discrepancies explicitly, thus exhibiting degraded performance under severe graph domain skew.
@@ 3 Challenges and Insights
Representation divergence manifests as high-variance, often conflicting gradients. Baseline FedAvg oscillates on a rugged loss surface (Fig. 2(b)), while a state-of-the-art aggregator, FGGP, remains unstable across settings (Fig. 2(c)). Simply exchanging larger messages does not tame the noise, such as prototypes, local statistics or second moments; the optimiser continues to chase contradictory directions.